OpenStudy (anonymous):

Solve. 3x ≤ 21 also can someone explain to me what this ≤ means.It means equal.

OpenStudy (anonymous):

\[3x \leq 21 \] \[x \leq 21/3 \] \[x \leq 7 \] The \(\leq\) means less than or equal to so the answer is "x is less than or equal to 7"